# Bounds on the number of squares in recurrence sequences

### v2

@@ 1.2. Notation
[…]
@@ 1.3. Conjectures
If $-N_{\alpha}$ is a perfect square, then there are at most two distinct integer squares among the $y_{k}$’s.
@@ 1.4. Results
In this paper, we prove Conjecture 1.2 when $-N_{\alpha}$ is a square, except for a very limited set of sequences, where we prove a slightly weaker result.
@@
Let $b=1$, $a$ and $d$ be positive integers, where $d$ is not a square, $N_{\alpha}<0$ and $-N_{\alpha}$ is a square.
@@ 2. Diophantine Approximation via Hypergeometric Functions
For any algebraic integers $p$ and $q$ in $\mathbb{K}$, let $r_{0}$ be the smallest positive integer such that $\left(Q-1/E\right)\ell_{0}|q|/\left(Q-1\right)<cE^{r_{0}}$, where $10<c<$ 1.
@@ 2.1. Construction of Approximations
Let $t^{\prime}$, $u_{1}$ and $u_{2}$ be rational integers with $t^{\prime}<0$  [^1] To avoid confusion with $t$ in the expression for $\varepsilon$, we use $t^{\prime}$ here to denote what was $t$ in our previous works like [16]. Similarly, we will use $d^{\prime}$ where $d$ was used in previous works.. We let $u=\left(u_{1}+u_{2}\sqrt{t^{\prime}}\right)/2$ be an algebraic integer in $\mathbb{K}=\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$ with $\sigma(u)=\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/2$ as its algebraic (and complex) conjugate. Put $\omega=u/\sigma(u)$ and write $\omega=e^{i\varphi}$, where $-\pi<\varphi\leq\pi$. For any real number $\nu$, we shall put $\omega^{\nu}=e^{i\nu\varphi}$ – unless otherwise stated, we will use this convention throughout this paper.
@@
$X_{m,n,r}(z)={}_{2}F_{1}(-r-\nu,-r,1-\nu,z),\quad Y_{m,n,r}=z^{r}X_{m,n,r}\left(z^{-1}\right)$
@@
(2.8) $g_{3}$ =$=\left\{\begin{array}[]{ll}1&\text{if $t^{\prime}\equiv 1\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 2&\text{if $t^{\prime}\equiv 3\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 4&\text{otherwise,}\end{array}\right.$
@@
Since $d^{\prime}=u_{2}^{2}t^{\prime}$ and $\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/(2g)$ is an algebraic integer, it follows that $p_{r}$ is an algebraic integer in $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$. The analogous expression for $q_{r}$ shows that it is also an algebraic integer.
@@
$\ell_{0}=\mathcal{C}_{4,2}|\varphi|=0.2|\varphi|.$
@@ 3.1. Representation Proposition
Let $a\neq 0$,  [^2] PV: why do I want $a\neq 0$? $a=0$ will give me the squares among the $x_{k}$’s from $\varepsilon^{2k}$. $b>0$ and $d$ be rational integers such that $d$ is not a square. Put $\alpha=a+b^{2}\sqrt{d}$ and denote $N_{\mathbb{Q}(\sqrt{d})/\mathbb{Q}}(\alpha)$ by $N_{\alpha}$. Suppose that $N_{\alpha}$ is not a square, $x\neq 0$ and $y>0$ are rational integers with
@@
(b) If $-N_{\alpha}$ is a square, then $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, where the first relationship in part (a) is replaced by
±$\pm f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}.$
@@
(3.3) $v_{p}\left(g_{1}^{2}\right)=\left\{\begin{array}[]{ll}2k_{1}+k_{2},&\text{if $k_{2}$ is even and $2k_{1}\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is even and $2k_{1}\geq k_{2}$,}\\ 2k_{1}+k_{2}-1,&\text{if $k_{2}$ is odd and $2k_{1}-1\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is odd and $2k_{1}+1\geq k_{2}$.}\end{array}\right.$
@@
(3.5) $v_{2}\left(g_{1}^{2}\right)=\left\{\begin{array}[]{ll}2k_{1}+k_{2}+2,&\text{if$k_{2}$ is even and $2k_{1}+2\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is even and $2k_{1}+2\geq k_{2}$,}\\ 2k_{1}+k_{2}+1,&\text{if $k_{2}$ is odd and $2k_{1}+1\leq k_{2}$,}\\ 2k_{2},&\text{if $k_{2}$ is odd and $2k_{1}+1\geq k_{2}$.}\end{array}\right.$
@@
$2\displaystyle 2v_{2}(y)+$ 2 $\geq\min\left(v_{2}(2atu),v_{2}\left(b^{2}\left(4N_{\varepsilon}+2du^{2}\right)\right)\right)$
@@
Lastly, we choose which of the two possible values of $r_{1}$ to use. As in the statement of Lemma 3.3, we let $r_{1}^{\prime}=tb^{2}+au+2by$ and $r_{1}^{\prime\prime}=tb^{2}+au-2by$ be the two possibilities. We will use Lemma 3.3(b). We divide the prime divisors of $\operatorname{sf}\left(N_{\alpha}\right)$ into three disjoint sets:
$I^{\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\},$
$I^{\prime\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)<v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\right\},$
$I^{\prime\prime\prime}$ =$=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)=v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\}.$
@@
We can write $(r+si)/(1+i)=(r+s)/2+(s-r)i/2$. Since $r$ and $s$ are both odd and $(1+i)^{4}=-4$, we find that $(s\pm r)/2$ are both integers and
$-\left(\frac{2b^{2}r}{g_{1}^{2}}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left((r+s)/2+(s-r)i/2\right)^{4}.$
@@
Since $-N_{\alpha}$ is a square, we have $N_{\alpha}\not\equiv 1\pmod{4}$, so by Lemmas 3.4 and 3.3(c), we see that $\left(4b^{2}r/g_{1}^{2}\right)|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$. Hence $2b^{2}r/g_{1}^{2}$ is a divisor of $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ and so we take $f=2b^{2}r/g_{1}^{2}$ in this case.
@@
From (3.4), we have $4y^{2}=2atu+2b^{2}du^{2}+4b^{2}N_{\varepsilon}$. So $atu+b^{2}du^{2}$ must be even. I.e., $atu$ and $b^{2}du^{2}$ have the same parity.
@@
If $d\equiv 1\pmod{8}$, then $4N_{\varepsilon}=t^{2}-du^{2}\equiv 0\pmod{8}$ and so $N_{\varepsilon}$ is even. Furthermore, since $abd$ is odd, we have $N_{\alpha}$ is even, so $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ is also even and the desired conclusion holds.
@@ 3.2. Lower bounds for $y_{k}$’s
[…]
@@
In fact, if $(d,t,u)\neq(5,1,1)$, then we can replace $52du^{2}/$ 5 by $85du^{2}/$ 8 and if $N_{\varepsilon}=1$, then we can replace $52du^{2}/$ 5 by $du^{2}$.
@@
From the second equality in (3.14), equation (3.13) follows for $k=1$.
Writing $\varepsilon^{\ell}=t_{\ell}+u_{\ell}\sqrt{d}$ for $\ell\geq 1$ (note that $u_{1}=u$ in the notation of this lemma), we can show by an easy induction that $\left(u_{\ell}\right)_{\ell\geq 1}$ is an increasing sequence. Hence, for $k<0$,
@@
We will also need to know when $K$ in Lemma 3.5 is not equal to $-1$ for $b=1$.
@@
[…]
@@
In these cases, we have $y_{-2}=a^{2}+1$ and $y_{-2}=4a^{2}+1$, respectively. So in each of these cases, $K=-2$.
@@
$a^{2}-t^{2}=-8$ can only occur for $a=1$ and $t=3$ (since we are assuming that $a$ and $t$ are positive integers). But in this case, $b^{2}t-au=2$, so $84y_{-1}=\left(b^{2}t-au\right)^{2}-N_{\alpha}u^{2}=$ 8.
@@
In case (2), we have $d=\left(t^{2}\pm 4\right)/4$ and $a^{2}-d=-1$. We proceed similarly, obtaining $84a^{2}-t^{2}=-$ 8 (which is never possible) when $N_{\varepsilon}=1$ and $04a^{2}-t^{2}=$ 0 when $N_{\varepsilon}=-1$. Here we have $t=2a$ and $d=a^{2}+1$, so $\alpha=\varepsilon$. Once again, $b^{2}t-au=0$, so $y_{-1}=1$.
@@ 3.3. Gap Principle
Let $\omega=e^{i\theta}$ with $-\pi<\theta\leq\pi$.
@@
for $-\pi<\varphi\leq\pi$ such that $e^{i\varphi}=e^{-i\theta/4}z$. Squaring this quantity and simplifying, we obtain
@@
Since $\cos(4\varphi)=8\cos^{4}(\varphi)-8\cos^{2}(\varphi)+1$, we have $016\cos^{4}(\varphi)-16\cos^{2}(\varphi)+c_{0}^{2}=$ 0 and using this relation, we find that $c_{0}^{2}/\left(8\cos^{2}(\varphi)(\cos(\varphi)+1)\right)$ is a root of the polynomial $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$. Part (b) now follows and follows with inequalities too, since the smallest positive real root of $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$ grows with $c_{0}$. ∎
@@
[…]
(a) Suppose that $-N_{\alpha}$ is a square. If $y_{i}$ and $y_{j}$ are distinct squares with $i,j\neq 0$ and $y_{j}>y_{i}\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},b^{2}\left|N_{\alpha}\right|/d\right\}$, then
@@
(b) Suppose $-N_{\alpha}$ is not a square. If $y_{k_{1}}$, $y_{k_{2}}$ and $y_{k_{3}}$ are three distinct squares with $k_{1},k_{2},k_{3}\neq 0$ and $y_{k_{3}}>y_{k_{2}}>y_{k_{1}}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, then there exist distinct $i,j\in\{k_{1},k_{2},k_{3}\}$ such that
@@
If $y_{j}>y_{i}\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},4.27b^{2}\left|N_{\alpha}\right|^{2}/d\right\}$, then we can replace $15.36$ with $182$.
@@
±$\pm f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\quad\text{and}$
@@
where $f_{j}$ satisfies $f_{j}|\left(4b^{2}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)\right)$ and $f_{j}^{2}<16b^{4}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$. Recall that the other terms in the relationships for $f_{j}$ in Proposition 3.1 are not present here since $N_{\varepsilon}=\pm 1$.
For any two distinct squares among the $y_{k}$’s, say $y_{i}$ and $y_{j}$ with $i\neq 0$ and $j\neq 0$, we will assume that the $\pm$ on the left-hand side of (3.3) is always $+$. That is,
(3.17) $f_{i}^{2}\left(x_{i}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\text{ and }$
$f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4},$
@@
$=$ $2i\operatorname{Im}\left(f_{j}^{2}\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\right)=\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|},$
@@
(3.19) $\left|\omega-\left(\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$ =|±$=\left|\frac{\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|}}{\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}}\right|$
@@
$\left(x_{i}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)y_{j}=\pm\left(x_{j}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)y_{i}.$
Comparing the imaginary parts of both sides of this equation, we find that $y_{i}=y_{j}$, but this contradicts our assumption that $y_{j}>y_{i}$. Hence the left-hand side of (3.22) cannot be 00.
@@
(a) Suppose that $-N_{\alpha}$ is a square. From (3.24), we see that at least one of $1\pm\sqrt{-1}$ always divides the numerator of the right-hand side of (3.23) and so
@@
Furthermore, from Proposition 3.1(b), we know that $f_{i},f_{j}|b^{2}$, since $-N_{\alpha}$ is a square. So
@@
(b) We again use (3.23) and (3.24). Notice that $x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}$ on the left-hand side of (3.24) for $\zeta_{4}^{(j)}=\pm\sqrt{-1}\zeta_{4}^{(i)}$ can be as small as $1/\left(2\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}y\right)$ when $-N_{\alpha}$ is not a square, so the proof of part (a) breaks down here.
Among the three values $\zeta_{4}^{(k_{1})}$, $\zeta_{4}^{(k_{2})}$ and $\zeta_{4}^{(k_{3})}$, there must be at least one pair such that one member of the pair is ±$\pm 1$ times the other. Choose any such pair and let $i$ and $j$ be the associated indices, ordered so that $y_{i}<y_{j}$.

### v3

@@ 1.2. Notation
(1.2) $y_{1}=\frac{b^{2}\left(t^{2}+du^{2}\right)+2atu}{4},\hskip 8.53581pty_{-1}=\frac{b^{2}\left(t^{2}+du^{2}\right)-2atu}{4}$
@@ 1.3. Conjectures
If $N_{\alpha}$ is a perfect square, then there are at most two distinct integer squares among the $y_{k}$’s.
@@ 1.4. Results
In this paper, we prove Conjecture 1.2 when $N_{\alpha}$ is a square, except for a very limited set of sequences, where we prove a slightly weaker result.
@@
Let $b=1$, $a$ and $d$ be positive integers, where $d$ is not a square, $N_{\alpha}<0$ and $N_{\alpha}$ is a square.
@@ 2. Diophantine Approximation via Hypergeometric Functions
For any algebraic integers $p$ and $q$ in $\mathbb{K}$, let $r_{0}$ be the smallest positive integer such that $\left(Q-1/E\right)\ell_{0}|q|/\left(Q-1\right)<cE^{r_{0}}$, where $0<c<1$.
@@ 2.1. Construction of Approximations
Let $t^{\prime}$, $u_{1}$ and $u_{2}$ be rational integers with $t^{\prime}<0$  [^1] To avoid confusion with $t$ in the expression for $\varepsilon$, we use $t^{\prime}$ here to denote what was $t$ in our previous works like [16]. Similarly, we will use $d^{\prime}$ where $d$ was used in previous works. such that $u=\left(u_{1}+u_{2}\sqrt{t^{\prime}}\right)/2$ be an algebraic integer in $\mathbb{K}=\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$ with $\sigma(u)=\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/2$ as its algebraic (and complex) conjugate. Put $\omega=u/\sigma(u)$ and write $\omega=e^{i\varphi}$, where $\pi<\varphi\leq\pi$. For any real number $\nu$, we shall put $\omega^{\nu}=e^{i\nu\varphi}$ – unless otherwise stated, we will use this convention throughout this paper.
@@
$X_{m,n,r}(z)={}_{2}F_{1}(-r-\nu,-r,1-\nu,z),\quad Y_{m,n,r}(z)=z^{r}X_{m,n,r}\left(z^{-1}\right)$
@@
(2.8) $g_{3}$ $=\left\{\begin{array}[]{ll}1&\text{if $t^{\prime}\equiv 1\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 2&\text{if $t^{\prime}\equiv 3\pmod{4}$ and $\left(u_{1}-u_{2}\right)/g_{1}\equiv 0\pmod{2}$},\\ 4&\text{otherwise,}\end{array}\right$.
@@
Since $d^{\prime}=u_{2}^{2}t^{\prime}/g^{2}$ and $\left(u_{1}-u_{2}\sqrt{t^{\prime}}\right)/(2g)$ is an algebraic integer, it follows that $p_{r}$ is an algebraic integer in $\mathbb{Q}\left(\sqrt{t^{\prime}}\right)$. The analogous expression for $q_{r}$ shows that it is also an algebraic integer.
@@
(2.14) $\ell_{0}=\mathcal{C}_{4,2}|\varphi|=0.2|\varphi|.$
@@ 3.1. Representation Proposition
Let $a\neq 0$, $b>0$ and $d$ be rational integers such that $d$ is not a square. Put $\alpha=a+b^{2}\sqrt{d}$ and denote $N_{\mathbb{Q}(\sqrt{d})/\mathbb{Q}}(\alpha)$ by $N_{\alpha}$. Suppose that $N_{\alpha}$ is not a square, $x\neq 0$ and $y>0$ are rational integers with
@@
(b) If $N_{\alpha}$ is a square, then $f|\left(b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$, where the first relationship in part (a) is replaced by
$\pm f^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left(r+s\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}.$
@@
[…]
@@
[…]
@@
$2v_{2}(y)+2$ $\geq\min\left(v_{2}(2atu),v_{2}\left(b^{2}\left(4N_{\varepsilon}+2du^{2}\right)\right)\right)$
@@
Lastly, we choose which of the two possible values of $r_{1}$ to use. As in the statement of Lemma 3.3, we let $r_{1}^{\prime}=tb^{2}+au+2by$ and $r_{1}^{\prime\prime}=tb^{2}+au-2by$ be the two possibilities. We let $g_{1}^{\prime}$ and $g_{1}^{\prime\prime}$ be the associated values of $g_{1}$. We will use Lemma 3.3(b). We divide the prime divisors of $\operatorname{sf}\left(N_{\alpha}\right)$ into three disjoint sets:
$I^{\prime}$ $=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)<v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\},$
$I^{\prime\prime}$ $=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)<v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)\right\},$
$I^{\prime\prime\prime}$ $=\left\{p:p\text{ prime, with }p|\operatorname{sf}\left(N_{\alpha}\right)\text{ and }v_{p}\left(r_{1}^{\prime}/g_{1}^{\prime 2}\right)=v_{p}\left(r_{1}^{\prime\prime}/g_{1}^{\prime\prime 2}\right)\right\}.$
@@
We can write (r+si)/(1+i)=(r+s)/2+(sr)i/2(r+si)/(1+i)=(r+s)/2+(s-r)i/2. Since $r$ and $s$ are both odd and $(1+i)^{4}=-4$, we find that $(s\pm r)/2$ are both integers and
$\left(\frac{2b^{2}r}{g_{1}^{2}}\right)^{2}\left(x+N_{\varepsilon}\sqrt{N_{\alpha}}\right)=\left(a+\sqrt{N_{\alpha}}\right)\left((r+s)/2+(s-r)i/2\right)^{4}.$
@@
Since $N_{\alpha}$ is a square, we have $N_{\alpha}\not\equiv 1\pmod{4}$, so by Lemmas 3.4 and 3.3(c), we see that $\left(4b^{2}r/g_{1}^{2}\right)|\left(2b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)\right)$. Hence $2b^{2}r/g_{1}^{2}$ is a divisor of $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ and so we take $f=2b^{2}r/g_{1}^{2}$ in this case.
@@
From (3.4), we have $4y^{2}=2atu+2b^{2}du^{2}+4b^{2}N_{\varepsilon}$. So $atu+b^{2}du^{2}$ must be even. I.e., atuatu and $b^{2}du^{2}$ have the same parity.
@@
If $d\equiv 1\pmod{8}$, then $4N_{\varepsilon}=t^{2}-du^{2}\equiv 0\pmod{8}$ and so $N_{\varepsilon}$ is even. Furthermore, since abdabd is odd, we have $N_{\alpha}$ is even, so $b^{2}\operatorname{rad}\left(\gcd\left(uN_{\alpha},N_{\varepsilon}\right)\right)$ is also even and the desired conclusion holds.
@@ 3.2. Lower bounds for $y_{k}$’s
Let the $y_{k}$’s be defined by (1.1) with the notation and assumptions there. Suppose that $N_{\alpha}<0$. Let $K$ be the largest negative integer such that $y_{K}>b^{2}$.
@@
In fact, if $(d,t,u)\neq(5,1,1)$, then we can replace $2du^{2}/5$ by $5du^{2}/8$ and if $N_{\varepsilon}=1$, then we can replace $2du^{2}/5$ by $du^{2}$.
@@
From the second equality in (3.14) and $N_{\alpha}<0$, equation (3.13) follows for $k=1$.
Writing $\varepsilon^{\ell}=t_{\ell}+u_{\ell}\sqrt{d}$ for $\ell\geq 1$ (note that $u_{1}=u/2$ in the notation of this lemma), we can show by an easy induction that $\left(u_{\ell}\right)_{\ell\geq 1}$ is an increasing sequence. Hence, for $k<0$,
@@
We will also need to know when $K$ in Lemma 3.5 is not equal to 1-1 for $b=1$.
@@
Let the $y_{k}$’s be defined by (1.1) with the notation and assumptions there. Suppose that $b=1$ and $N_{\alpha}<0$.
@@
In these cases, we have $y_{-2}=a^{2}+1$ and $y_{-2}=4a^{2}+1$, respectively. So in each of these cases, K=2K=-2.
@@
$a^{2}-t^{2}=-8$ can only occur for $a=1$ and $t=3$ (since we are assuming that $a$ and $t$ are positive integers). But in this case, $b^{2}t-au=2$, so $4y_{-1}=\left(b^{2}t-au\right)^{2}-N_{\alpha}u^{2}=8$.
@@
In case (2), we have $d=\left(t^{2}\pm 4\right)/4$ and $a^{2}-d=-1$. We proceed similarly, obtaining $4a^{2}-t^{2}=-8$ (which is never possible) when $N_{\varepsilon}=1$ and $4a^{2}-t^{2}=0$ when $N_{\varepsilon}=-1$. Here we have $t=2a$ and $d=a^{2}+1$, so $\alpha=\varepsilon$. Once again, $b^{2}t-au=0$, so $y_{-1}=1$.
@@ 3.3. Gap Principle
Let $\omega=e^{i\theta}$ with $\pi<\theta\leq\pi$.
@@
for $\pi<\varphi\leq\pi$ such that $e^{i\varphi}=e^{-i\theta/4}z$. Squaring this quantity and simplifying, we obtain
@@
Since $\cos(4\varphi)=8\cos^{4}(\varphi)-8\cos^{2}(\varphi)+1$, we have $16\cos^{4}(\varphi)-16\cos^{2}(\varphi)+c_{0}^{2}=0$ and using this relation, we find that $c_{0}^{2}/\left(8\cos^{2}(\varphi)(\cos(\varphi)+1)\right)$ is a root of the polynomial $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$. Part (b) now follows and follows with inequalities too, since the smallest positive real root of $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$ grows with $c_{0}$. ∎
@@
Let the $y_{k}$’s be defined as in (1.1) with $N_{\alpha}<0$.
(a) Suppose that $N_{\alpha}$ is a square. If $y_{i}$ and $y_{j}$ are distinct squares with $i,j\neq 0$ and $y_{j}>y_{i}\geq\max\left(4\sqrt{\left|N_{\alpha}\right|/d},b^{2}\left|N_{\alpha}\right|/d\right)$, then
@@
(b) Suppose $N_{\alpha}$ is not a square. If $y_{k_{1}}$, $y_{k_{2}}$ and $y_{k_{3}}$ are three distinct squares with $k_{1},k_{2},k_{3}\neq 0$ and $y_{k_{3}}>y_{k_{2}}>y_{k_{1}}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, then there exist distinct $i,j\in\{k_{1},k_{2},k_{3}\}$ such that
@@
If $y_{j}>y_{i}\geq\max\left(4\sqrt{\left|N_{\alpha}\right|/d},4.27b^{2}\left|N_{\alpha}\right|^{2}/d\right)$, then we can replace $15.36$ with $182$.
@@
$\pm f_{j}^{2}\left(x_{j}+N_{\varepsilon^{j}}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\quad\text{and}$
@@
where $f_{j}$ satisfies $f_{j}|\left(4b^{2}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)\right)$ and $f_{j}^{2}<16b^{4}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$. Recall that the other terms in the relationships for $f_{j}$ in Proposition 3.1 are not present here since $N_{\varepsilon^{j}}=\pm 1$.
For any two distinct squares among the $y_{k}$’s, say $y_{i}$ and $y_{j}$ with $i\neq 0$ and $j\neq 0$, we will assume that the $\pm$ on the left-hand side of (3.3) is always ++. That is,
(3.17) $f_{i}^{2}\left(x_{i}+N_{\varepsilon^{i}}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\text{ and }$
$f_{j}^{2}\left(x_{j}+N_{\varepsilon^{j}}\sqrt{N_{\alpha}}\right)$ $=\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4},$
@@
$=$ $2i\operatorname{Im}\left(f_{j}^{2}\left(x_{j}+N_{\varepsilon^{j}}\sqrt{N_{\alpha}}\right)\right)=\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|},$
@@
(3.19) $\left|\omega-\left(\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$ $=\left|\frac{\pm 2f_{j}^{2}\sqrt{\left|N_{\alpha}\right|}}{\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}}\right|$
@@
$\left(x_{i}+N_{\varepsilon^{i}}\sqrt{N_{\alpha}}\right)y_{j}=\pm\left(x_{j}+N_{\varepsilon^{j}}\sqrt{N_{\alpha}}\right)y_{i}.$
Comparing the imaginary parts of both sides of this equation, we find that $y_{i}=y_{j}$, but this contradicts our assumption that $y_{j}>y_{i}$. Hence the left-hand side of (3.22) cannot be 0.
@@
(a) Suppose that $N_{\alpha}$ is a square. From (3.24), we see that at least one of $1\pm\sqrt{-1}$ always divides the numerator of the right-hand side of (3.23) and so
@@
Furthermore, from Proposition 3.1(b), we know that $f_{i},f_{j}|b^{2}$, since $N_{\alpha}$ is a square. So
@@
(b) We again use (3.23) and (3.24). Notice that $x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}$ on the left-hand side of (3.24) for $\zeta_{4}^{(j)}=\pm\sqrt{-1}\zeta_{4}^{(i)}$ can be as small as $1/\left(2\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}y\right)$ when $N_{\alpha}$ is not a square, so the proof of part (a) breaks down here.
Among the three values $\zeta_{4}^{(k_{1})}$, $\zeta_{4}^{(k_{2})}$ and $\zeta_{4}^{(k_{3})}$, there must be at least one pair such that one member of the pair is $\pm 1$ times the other. Choose any such pair and let $i$ and $j$ be the associated indices, ordered so that $y_{i}<y_{j}$.